OpenStudy (anonymous):

a. ultraviolet rays, which cause sun burn by penetrating only the top level of skin, have a wavelength of ultraviolet rays.show your work. b. the wavelengths of visible light are between 3.8x10^-7 m and 7.6x10^-7m. are these wavelength longer or shorter than those of ultraviolet rays? explain I have figured out the first one but i cant figure out the second one! please help

OpenStudy (anonymous):

If you know the wavelength of ultraviolet rays you can just look at which is bigger numerically. If you have difficulty discerning this on sight you can divide the wavelength of light by that of ultraviolet rays. \[= 1 \] then they have the same wavelength if \[< 1\] then the wavelength of visible light is shorter than that of ultraviolet and if \[>1\] then the wavelength of ultraviolet let is shorter than that of visible and if

OpenStudy (agent0smith):

When you compare numbers that have the same exponent (the power on the 10) in scientific notation , just compare the two numbers in front of the 10... eg \[3.8 \times 10 ^{-7}\] and \[7.6 \times 10 ^{-7} \] have the same exponent, so just compare 3.8 and 7.6. 7.6 is bigger, so it has a longer wavelength. When they have different powers on the 10, then just compare the exponents... eg \[3.8 \times 10 ^{-7}\] and \[7.6 \times 10 ^{-8} \] 10^-8 is *smaller* than 10^-7, so it has the shorter wavelength.